# Baichuan Alignment Technical Report

### v2

@@ Abstract
We introduce Baichuan Alignment, a detailed analysis of the alignment techniques employed in the Baichuan series of models. This represents the industry’s first comprehensive account of alignment methodologies, offering valuable insights for advancing AI research. We investigate the critical components that enhance model performance during the alignment process, including optimization methods, data strategies, capability enhancements, and evaluation processes. The process spans three key stages: Prompt Augmentation System (PAS), Supervised Fine-Tuning (SFT), and Preference Alignment. The problems encountered, the solutions applied, and the improvements made are thoroughly recorded.
Through comparisons across well-established benchmarks, we highlight the technological advancements enabled by Baichuan Alignment. Baichuan-Instruct is an internal model, while Qwen2-Nova-72B and Llama3-PBM-Nova-70B are instruct versions of the Qwen2-72B and Llama-3-70B base models, optimized through Baichuan Alignment. Baichuan-Instruct demonstrates significant improvements in core capabilities, with user experience gains ranging from 17% to 28%, and performs exceptionally well on specialized benchmarks. In open-source benchmark evaluations, both Qwen2-Nova-72B and Llama3-PBM-Nova-70B consistently outperform their respective official instruct versions across nearly all datasets. This report aims to clarify the key technologies behind the alignment process, fostering a deeper understanding within the community. Llama3-PBM-Nova-70B model is available at https://huggingface.co/PKU-Baichuan-MLSystemLab/Llama3-PBM-Nova-70B.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have achieved significant breakthroughs [67, 23, 69, 10, 18, 1, 3, 4, 81, 79], showing early signs of advancing toward Artificial General Intelligence (AGI). Through the mechanism of next-token prediction, LLMs undergo self-supervised pre-training on vast datasets, thereby acquiring a diverse spectrum of capabilities that empower them to perform an array of tasks, such as text continuation, summarization, and creative composition. Further advancements in alignment methodologies, such as Prompt Augmentation Systems (PAS) [103], Supervised Fine-Tuning (SFT), and preference modeling [65], have become important to the evolution of LLMs. These developments have enhanced the models’ ability to comprehend user intentions and adhere to instructions, thereby enhancing their conversational skills and adaptability to complex real-world scenarios. Despite its critical importance, a comprehensive understanding of alignment remains largely inaccessible to the broader public [5, 95, 84, 19, 98, 8, 55, 31, 37, 94, 81, 1]. The broad and intricate nature of alignment results in fragmented research efforts, which often provide only a narrow insight into specialized areas, making it challenging to present a holistic perspective of the alignment landscape. Additionally, alignment is frequently regarded as a proprietary cornerstone in the LLMs training, shrouded in corporate confidentiality. To foster advancement within the LLM community, we present a comprehensive and systematic exposition of Baichuan Alignment, featuring a suite of advanced and practical alignment techniques.
@@
Baichuan Alignment comprises three critical phases: Prompt Augmentation Systems (PAS), Supervised Fine-Tuning (SFT) and Preference Alignment. The PAS stage aims to transform user queries into instructions that are more comprehensible and actionable for LLMs through automated Prompt Engineering (PE) techniques. The SFT stage equips LLMs with the ability to engage in dialogue and handle complex tasks using a large corpus of high-quality and diverse data. The Preference Alignment further aligns the LLMs with human values and preferences. This report primarily focuses on four aspects: optimization, data, key capability enhancement, and system evaluation, which are critical elements of Baichuan Alignment. The optimization ensures the effectiveness and efficiency of LLM training, and Section 2 provides a detailed discussion on methods for training, prompt argumentation, and model merging, which collectively accelerate and improve model performance. In Section 3, we emphasize the importance of data in alignment, with a focus on prompt selection, response construction, and preference data. Section 4 outlines the challenges encountered in enhancing core capabilities on the path to AGI, detailing the specific approaches and insights gained through Baichuan Alignment. Finally, Section 5 presents a comprehensive evaluation of Baichuan Alignment, focusing on user experience, open-source benchmarks, and specific capability assessments. This evaluation is crucial for assessing model capabilities and guiding iterative improvements, particularly through user evaluations and third-party product assessments that align closely with user experience. Additionally, we introduce tailored benchmarks, including CFBench [99], SysBench [66], and FB-Bench [52], designed to address current challenges and practical needs in LLM applications.
@@
We conducted a systematic evaluation of multiple models after Baichuan Alignment from various perspectives. Among these, Baichuan-Instruct serves as our internal model, while Qwen2-Nova-72B and Llama3-PBM-Nova-70B are instruct versions derived from the Qwen2-72B [95] and Llama-3-70B [27] base models, respectively, fine-tuned using Baichuan Alignment techniques. User experience assessments reveal that Baichuan-Instruct exhibits significant improvements across several core capabilities, with enhancements ranging from 17% to 28%. Notably, there are marked improvements in mathematics and reasoning, with increases of 28% and 23%, respectively. In terms of open-source benchmarks, Qwen2-Nova-72B demonstrates substantial advancements over its official instruct version, Qwen2-72B-Instruct, across multiple leaderboards. Similarly, Llama3-PBM-Nova-70B shows significant improvements compared to Llama-3-70B-Instruct, particularly achieving a 60% relative increase (from 46.6 to 74.5) on the ArenaHard benchmark. On specific benchmarks, Baichuan-Instruct ranks competitively in core competencies on CFBench, SysBench, and FBBench, compared to the leading models currently available. This multidimensional, comprehensive, and in-depth evaluation underscores the superiority of Baichuan Alignment technology and its applicability across different foundational models. As a crucial step toward AGI, we are publicly sharing the challenges encountered during the Baichuan Alignment process, the solutions devised, and some in-depth insights. This disclosure aims to provide the community with new perspectives or lessons learned from failures, thereby fostering discussions on alignment and making a meaningful contribution to the advancement toward AGI.
@@ 2.1 Training
where the is a adjustable coefficient, and are the normalized annotated absolute score for the chosen and rejected data. Note that in our preference dataset, we also mixed some dataset without absolute score, like open source hh-rlhf[7] dataset, SHP[28] dataset. For data in these dataset, the is set to 0. With this objective, we find the fitted reward model is more robust. The RM model is trained for 1 to 2 epochs, depending on its size.
@@
In our experiment, we found that even without the critic model, GRPO can still achieve comparable results with PPO. The evaluation benchmark(described in Section 5.2.1) difference between GRPO and PPO is for FollowBench, for CFBench PSR Full and for SysBench. On the other hand, GRPO can save nearly half of the training resources compared to PPO. Additionally, it achieves better performance than direct optimization methods like DPO[68] and KTO[29], while requiring only a marginal increase in training resources. Therefore, the GRPO is chosen as our Reinforcement Learning method.
@@ 3 Data
Alignment data has been extensively validated as critical to the ultimate performance of LLMs, includeing both prompt response pairs and preference data [46, 90, 44, 26, 30]. Figure 4 delineates the comprehensive pipeline for constructing the Baichuan alignment dataset. Initially, we develop an prompt system and a classification model (Section 3.1.1), which form the foundational basis of the data flywheel. We then elaborate on the pivotal steps in data construction, focusing on prompt diversity (Section 3.1.2) and prompt quality (Section 3.1.3). Furthermore, we detail various techniques and best practices in generating responses (Section 3.2). Additionally, we provide insights derived from the process of constructing preference data (Section 3.3).
@@ 4.1 Instruction Following
The ability to following instruction is a critical capability of any advanced Large Language Model (LLM), particularly when confronted with the intricacies of real-world applications. Baichuan Alignment has implemented several key enhancements to bolster this crucial capability, as depicted in Figure 5. These primarily include the construction of complex system messages, expansion of instruction constraints, response reversal, and textbook learning.
@@ 4.2 Math
In many datasets of math problem-solving, the provided answers often lack detailed steps. Manual annotation is a common method, however, some intermediate steps may be ignored. Nevertheless, these steps are important for model training. By leveraging large language models in combination with prompts and standard answers, we can generate detailed solutions that maintain a consistent style based on the reference answers. This approach synthesizes a substantial amount of math alignment data with detailed steps, significantly enhancing the model’s mathematical capabilities. Directly generated answers often struggle to match correct answers, so providing reference answers and steps helps the model enrich these into coherent reasoning chains. We find that reference answers significantly improve the quality of generated responses. However, since reference answers may contain errors, we instruct the model to consider them as having a 95% probability of correctness. This approach reduces incorrect inferences based on potentially faulty references. Furthermore, we perform post-processing to control the quality of the generated responses, for example, ensuring that they do not explicitly mention the existence of reference answers. Finally, the proportion of math data accounts for approximately 20% of the total data used for supervised fine-tuning.
@@ 4.4 Code
Code Knowledge Q&A is essentially common knowledge questioning and does not enhance coding skills, so it is downsampled by 80%. Data categorized as Very Difficult often requires knowledge beyond the model’s capabilities or extensive context in the prompt to solve, making it prone to hallucinations and difficult to verify accurately with human oversight, so this type of data is discarded. Additionally, we have balanced the data across different programming languages.
@@ 4.5 Tool-using
4.6 Prompt Augmentation System
We post-trained a 33B chat model with a curated dataset for the Prompt Augmentation System to generate prompt supplement content. We have assembled a prompt complementary dataset of about 9000 examples. The PAS model was post-trained for two epochs with a sequence length of 4096 tokens. Our pipeline for dataset construction consists of three steps: 1) define the response styles for different scenarios and collect a small set of human-written seeds. 2) few-shot a LLM for data synthesis. 3) improve data quality with LLM self-correct.
@@ 5 Evaluation
In this section, we will provide a detailed overview of the Baichuan alignment evaluation system, which, by incorporating user perception, open-source benchmarks, and a specially constructed evaluation set for key capabilities, effectively demonstrates the comprehensiveness and sufficiency of our evaluation, as well as the superiority of the alignment techniques. We assessed the leading models currently available, including GPT [1], Claude [4], Qwen [5, 95], ERNIE [77], Moonshot [61], Yi-Large [98], DeepSeek-V2 [8], GLM [31], mixtral [41] and Llama [81, 27].
@@ 5.1.2 Result and Discussion
Table 1 illustrates the comparative improvements in pass rate before and after the optimization actions mentioned in this paper, covering nearly all key tasks and capabilities of interest in Baichuan’s product iterations. Notably, represents the change in absolute percentage. Overall, the various optimization strategies discussed earlier have led to significant improvements across all tasks, with the most pronounced increases observed in Math (28%), KQA (25%), and Reason (23%). These improvements primarily originate from the optimizations discussed in Sections 4.2, 4.3, and 4.1. Notably, the enhancement in mathematical capabilities is significantly driven by data optimizations involving "comprehensive coverage of knowledge points and more detailed problem-solving steps," while the improvement in response quality is largely attributed to the implementation of "Reasoning CoT". The change demonstrated marked improvements in instruction following , attributed particularly to constraint expansion and response reversal engineering as discussed in the relevant Sections 4.1. The function call appears to exhibit the most gradual growth, primarily due to its inherent complexity and challenges. Nevertheless, the methodology involving refined high-quality annotated data confers distinct advantages, as detailed in Section 4.5. Role Play enhances performance predominantly through the construction of personalized data and preference alignment, as clearly articulated in Sections 3.2, 2.1 and 3.3. Notably, PAS (4.6), and preference alignment, along with the enhancement of data diversity and quality, collectively contribute to significant improvements across multiple core capabilities.
@@ 5.2.1 Benchmarks
We have curated a selection of widely recognized open-source benchmarks that cover a broad spectrum of capabilities to conduct a comprehensive and in-depth evaluation of models optimized through Baichuan alignment. The benchmarks BBH[78], MixEval-Hard[63], and Alpaca-Eval 2.0[50] are primarily utilized to assess general capabilities. Instruction-following abilities are evaluated using IFEval[105] and FollowBench[42], while conversational proficiency is assessed through ArenaHard[47] and MTBench[101]. HumanEval[14], MATH[36], and GPQA[71] specifically focus on evaluating coding, mathematical, and knowledge-based abilities, respectively.
@@ 5.2.2 Major Results and Discussion
Table 2 presents a comparative analysis of Qwen2-Nova-72B against other models across several authoritative open-source benchmarks. It is evident that Qwen2-Nova-72B significantly outperforms the official instruct version, Qwen2-72B-Instruct, derived from the Qwen2-72B base model across all evaluation sets. The most notable improvement is observed on the ArenaHard, where performance increased from 48.1 to 75.1. Substantial gains are also evident in BBH and Math. Furthermore, when compared to the most advanced LLMs currently available, Qwen2-Nova-72B achieves leading rankings on ArenaHard, MTBench, HumanEval, BBH, Math, and FollowBench. Table 3 illustrates the performance of the Llama3-PBM-Nova-70B model following Baichuan Alignment. Compared to the Llama-3-70B-Instruct model, which is based on the same foundational model, Llama3-PBM-Nova-70B demonstrates significant performance advantages on ArenaHard (74.5), MixEval (58.1), AlpacaEval2.0 (56.9), and GPQA (34). When compared to the most advanced LLMs, it ranks second and third on AlpacaEval2.0 and ArenaHard, respectively. Overall, the instruct models obtained through Baichuan Alignment on two different open-source base models consistently outperform their official instruct versions across multiple open-source benchmarks. This clearly demonstrates the versatility of Baichuan Alignment in enhancing foundational models and its comprehensive ability to improve performance across various benchmarks.

### v3

# Nova: A Practical and Advanced Alignment

@@ Abstract
We introduce Nova, a suite of practical alignment techniques employed in a series of empirically validated high-performing models. This represents the first comprehensive account of alignment methodologies, offering valuable insights for advancing AI research. We investigate the critical components that enhance model performance during the alignment process, including optimization methods, data strategies, capability enhancements, and evaluation processes. The process spans three key stages: Prompt Augmentation System (PAS), Supervised Fine-Tuning (SFT), and Preference Alignment. The problems encountered, the solutions applied, and the improvements made are thoroughly recorded.
Through comparisons across well-established benchmarks, we highlight the technological advancements enabled by Nova Alignment. Importantly, Qwen2-Nova-72B and Llama3-PBM-Nova-70B are instruct versions of the Qwen2-72B and Llama-3-70B base models, optimized through Nova. The Nova models show significant core improvements, with user experience gains of 17% to 28%, and excels on specialized benchmarks. In open-source benchmark evaluations, both Qwen2-Nova-72B and Llama3-PBM-Nova-70B consistently outperform their respective official instruct versions across nearly all datasets. This report aims to clarify the key technologies behind the alignment process, fostering a deeper understanding within the community. Llama3-PBM-Nova-70B model is available at https://huggingface.co/PKU-Baichuan-MLSystemLab/Llama3-PBM-Nova-70B.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have achieved significant breakthroughs [67, 23, 69, 10, 18, 1, 3, 4, 81, 79], showing early signs of advancing toward Artificial General Intelligence (AGI). Through the mechanism of next-token prediction, LLMs undergo self-supervised pre-training on vast datasets, thereby acquiring a diverse spectrum of capabilities that empower them to perform an array of tasks, such as text continuation, summarization, and creative composition. Further advancements in alignment methodologies, such as Prompt Augmentation Systems (PAS) [103], Supervised Fine-Tuning (SFT), and preference modeling [65], have become important to the evolution of LLMs. These developments have enhanced the models’ ability to comprehend user intentions and adhere to instructions, thereby enhancing their conversational skills and adaptability to complex real-world scenarios. Despite its critical importance, a comprehensive understanding of alignment remains largely inaccessible to the broader public [5, 95, 84, 19, 98, 8, 55, 31, 37, 94, 81, 1]. The broad and intricate nature of alignment results in fragmented research efforts, which often provide only a narrow insight into specialized areas, making it challenging to present a holistic perspective of the alignment landscape. Additionally, alignment is frequently regarded as a proprietary cornerstone in the LLMs training, shrouded in corporate confidentiality. To foster advancement within the LLM community, we present a comprehensive and systematic exposition of Nova Alignment, featuring a suite of advanced and practical alignment techniques.
@@
Nova Alignment comprises three critical phases: Prompt Augmentation Systems (PAS), Supervised Fine-Tuning (SFT) and Preference Alignment. The PAS stage aims to transform user queries into instructions that are more comprehensible and actionable for LLMs through automated Prompt Engineering (PE) techniques. The SFT stage equips LLMs with the ability to engage in dialogue and handle complex tasks using a large corpus of high-quality and diverse data. The Preference Alignment further aligns the LLMs with human values and preferences. This report primarily focuses on four aspects: optimization, data, key capability enhancement, and system evaluation, which are critical elements of Nova Alignment. The optimization ensures the effectiveness and efficiency of LLM training, and Section 2 provides a detailed discussion on methods for training, prompt argumentation, and model merging, which collectively accelerate and improve model performance. In Section 3, we emphasize the importance of data in alignment, with a focus on prompt selection, response construction, and preference data. Section 4 outlines the challenges encountered in enhancing core capabilities on the path to AGI, detailing the specific approaches and insights gained through Nova Alignment. Finally, Section 5 presents a comprehensive evaluation of Nova Alignment, focusing on user experience, open-source benchmarks, and specific capability assessments. This evaluation is crucial for assessing model capabilities and guiding iterative improvements, particularly through user evaluations and third-party product assessments that align closely with user experience. Additionally, we introduce tailored benchmarks, including CFBench [99], SysBench [66], and FB-Bench [52], designed to address current challenges and practical needs in LLM applications.
@@
We conducted a systematic evaluation of multiple models after Nova Alignment from various perspectives. Among these, Qwen2-Nova-72B and Llama3-PBM-Nova-70B are instruct versions derived from the Qwen2-72B [95] and Llama-3-70B [27] base models, respectively, fine-tuned using Nova Alignment techniques. User experience assessments reveal that Nova Alignment exhibits significant improvements across several core capabilities, with enhancements ranging from 17% to 28%. Notably, there are marked improvements in mathematics and reasoning, with increases of 28% and 23%, respectively. In terms of open-source benchmarks, Qwen2-Nova-72B demonstrates substantial advancements over its official instruct version, Qwen2-72B-Instruct, across multiple leaderboards. Similarly, Llama3-PBM-Nova-70B shows significant improvements compared to Llama-3-70B-Instruct, particularly achieving a 60% relative increase (from 46.6 to 74.5) on the ArenaHard benchmark. On specific benchmarks, Qwen2-Nova-72B ranks competitively in core competencies on CFBench, SysBench, and FBBench, compared to the leading models currently available. This multidimensional, comprehensive, and in-depth evaluation underscores the superiority of Nova Alignment technology and its applicability across different foundational models. As a crucial step toward AGI, we are publicly sharing the challenges encountered during the Nova Alignment process, the solutions devised, and some in-depth insights. This disclosure aims to provide the community with new perspectives or lessons learned from failures, thereby fostering discussions on alignment and making a meaningful contribution to the advancement toward AGI.
@@ 2.1 Training
where the is a adjustable coefficient, and are the normalized annotated absolute score for the chosen and rejected data. Note that in our preference dataset used by our open-source model, we also mixed some dataset without absolute score, like open source hh-rlhf[7] dataset, SHP[28] dataset. For data in these dataset, the is set to 0. With this objective, we find the fitted reward model is more robust. The RM model is trained for 1 to 2 epochs, depending on its size.
@@
In our experiment, we found that even without the critic model, GRPO can still achieve comparable results with PPO. The evaluation benchmark(described in Section 5.2.1) difference between GRPO and PPO is for FollowBench, for CFBench PSR Full. On the other hand, GRPO can save nearly half of the training resources compared to PPO. Additionally, it achieves better performance than direct optimization methods like DPO[68] and KTO[29], while requiring only a marginal increase in training resources. Therefore, the GRPO is chosen as our Reinforcement Learning method.
@@ 3 Data
Alignment data has been extensively validated as critical to the ultimate performance of LLMs, includeing both prompt response pairs and preference data [46, 90, 44, 26, 30]. Figure 4 delineates the comprehensive pipeline for constructing the Nova alignment dataset. Initially, we develop an prompt system and a classification model (Section 3.1.1), which form the foundational basis of the data flywheel. We then elaborate on the pivotal steps in data construction, focusing on prompt diversity (Section 3.1.2) and prompt quality (Section 3.1.3). Furthermore, we detail various techniques and best practices in generating responses (Section 3.2). Additionally, we provide insights derived from the process of constructing preference data (Section 3.3).
@@ 4.1 Instruction Following
The ability to following instruction is a critical capability of any advanced Large Language Model (LLM), particularly when confronted with the intricacies of real-world applications. Nova Alignment has implemented several key enhancements to bolster this crucial capability, as depicted in Figure 5. These primarily include the construction of complex system messages, expansion of instruction constraints, response reversal, and textbook learning.
@@ 4.2 Math
In many datasets of math problem-solving, the provided answers often lack detailed steps. Manual annotation is a common method, however, some intermediate steps may be ignored. Nevertheless, these steps are important for model training. By leveraging large language models in combination with prompts and standard answers, we can generate detailed solutions that maintain a consistent style based on the reference answers. This approach synthesizes a substantial amount of math alignment data with detailed steps, significantly enhancing the model’s mathematical capabilities. Directly generated answers often struggle to match correct answers, so providing reference answers and steps helps the model enrich these into coherent reasoning chains. We find that reference answers significantly improve the quality of generated responses. However, since reference answers may contain errors, we instruct the model to consider them as having a 95% probability of correctness. This approach reduces incorrect inferences based on potentially faulty references. Furthermore, we perform post-processing to control the quality of the generated responses, for example, ensuring that they do not explicitly mention the existence of reference answers.
@@ 4.4 Code
Code Knowledge Q&A is essentially common knowledge questioning and does not enhance coding skills, so it is downsampled. Data categorized as Very Difficult often requires knowledge beyond the model’s capabilities or extensive context in the prompt to solve, making it prone to hallucinations and difficult to verify accurately with human oversight, so this type of data is discarded. Additionally, we have balanced the data across different programming languages.
@@ 5 Evaluation
In this section, we will provide a detailed overview of the Nova alignment evaluation system, which, by incorporating user perception, open-source benchmarks, and a specially constructed evaluation set for key capabilities, effectively demonstrates the comprehensiveness and sufficiency of our evaluation, as well as the superiority of the alignment techniques. We assessed the leading models currently available, including GPT [1], Claude [4], Qwen [5, 95], ERNIE [77], Moonshot [61], Yi-Large [98], DeepSeek-V2 [8], GLM [31], mixtral [41] and Llama [81, 27].
@@ 5.1.2 Result and Discussion
Table 1 illustrates the comparative improvements in pass rate before and after the optimization actions mentioned in this paper, covering nearly all key tasks and capabilities of interest in product iterations. Notably, represents the change in absolute percentage. Overall, the various optimization strategies discussed earlier have led to significant improvements across all tasks, with the most pronounced increases observed in Math (28%), KQA (25%), and Reason (23%). These improvements primarily originate from the optimizations discussed in Sections 4.2, 4.3, and 4.1. Notably, the enhancement in mathematical capabilities is significantly driven by data optimizations involving "comprehensive coverage of knowledge points and more detailed problem-solving steps," while the improvement in response quality is largely attributed to the implementation of "Reasoning CoT". The change demonstrated marked improvements in instruction following , attributed particularly to constraint expansion and response reversal engineering as discussed in the relevant Sections 4.1. The function call appears to exhibit the most gradual growth, primarily due to its inherent complexity and challenges. Nevertheless, the methodology involving refined high-quality annotated data confers distinct advantages, as detailed in Section 4.5. Role Play enhances performance predominantly through the construction of personalized data and preference alignment, as clearly articulated in Sections 3.2, 2.1 and 3.3. Notably, PAS (2.3), and preference alignment, along with the enhancement of data diversity and quality, collectively contribute to significant improvements across multiple core capabilities.
@@ 5.2.1 Benchmarks
We have curated a selection of widely recognized open-source benchmarks that cover a broad spectrum of capabilities to conduct a comprehensive and in-depth evaluation of models optimized through Nova alignment. The benchmarks BBH[78], MixEval-Hard[63], and Alpaca-Eval 2.0[50] are primarily utilized to assess general capabilities. Instruction-following abilities are evaluated using IFEval[105] and FollowBench[42], while conversational proficiency is assessed through ArenaHard[47] and MTBench[101]. HumanEval[14], MATH[36], and GPQA[71] specifically focus on evaluating coding, mathematical, and knowledge-based abilities, respectively.
@@ 5.2.2 Major Results and Discussion
Table 2 presents a comparative analysis of Qwen2-Nova-72B against other models across several authoritative open-source benchmarks. It is evident that Qwen2-Nova-72B significantly outperforms the official instruct version, Qwen2-72B-Instruct, derived from the Qwen2-72B base model across all evaluation sets. The most notable improvement is observed on the ArenaHard, where performance increased from 48.1 to 75.1. Substantial gains are also evident in BBH and Math. Furthermore, when compared to the most advanced LLMs currently available, Qwen2-Nova-72B achieves leading rankings on ArenaHard, MTBench, HumanEval, BBH, Math, and FollowBench. Table 3 illustrates the performance of the Llama3-PBM-Nova-70B model following Nova Alignment. Compared to the Llama-3-70B-Instruct model, which is based on the same foundational model, Llama3-PBM-Nova-70B demonstrates significant performance advantages on ArenaHard (74.5), MixEval (58.1), AlpacaEval2.0 (56.9), and GPQA (34). When compared to the most advanced LLMs, it ranks second and third on AlpacaEval2.0 and ArenaHard, respectively. Overall, the instruct models obtained through Nova Alignment on two different open-source base models consistently outperform their official instruct versions across multiple open-source benchmarks. This clearly demonstrates the versatility of Nova Alignment in enhancing foundational models and its comprehensive ability to improve performance across various benchmarks.